Problem 1: Suppose is a random variable with mean 64.5 and variance 144. Use Chebyshev’s inequality to approximate the following probabilities.
1
Expert's answer
2019-11-25T12:00:25-0500

CHEBYSHEV’S INEQUALITY

Let "X" be a discrete rv with mean "\mu" and standard deviation "\sigma". Then, for any "k\geq1,"


"P(|X-\mu|\geq k\sigma)\leq{1 \over k^2}"

Given that "\mu=64.5, \sigma^2=144"

"a)\ \ P(44<X<85)"


"\sigma=\sqrt{144}=12"

"{85-44 \over 2}=20.5"

"k={20.5 \over 12}"

"P(44<X<85)=1-P(|X-\mu|\geq k\sigma)="

"=1-P\bigg(|X-64.5|\geq {20.5 \over 12}\cdot 12\bigg)\geq1-{1 \over (20.5/12)^2}"

"P(44<X<85)>0.6573"

"b) P(36<X<93)"


"\sigma=\sqrt{144}=12"

"{93-36 \over 2}=28.5"

"k={28.5 \over 12}"

"P(36<X<93)=1-P(|X-\mu|\geq k\sigma)="

"=1-P\bigg(|X-64.5|\geq {28.5 \over 12}\cdot 12\bigg)\geq1-{1 \over (28.5/12)^2}"

"P(36<X<93)>0.8227"